OpenStudy (anonymous):

∫[0,3,x^3+x,] antiderivative...the 3 is on top and the 0 bottom then x^3 + x.

OpenStudy (amistre64):

can we split up the argument? antiD them separatley?

OpenStudy (amistre64):

other than that, consider the rules:\[\frac{d}{dx}x^{n}\to\ n\ x^{n-1}\] and its reversal \[\int x^n\to\ \frac{x^{n+1}}{n+1}\]